OpenStudy (anonymous):

Metal screws weigh 3 kilograms per cubic meter. The screw are packaged in a cube whose edge is 1.5 meter. Find the weight of the cube. Round the answer to the nearest tenth of a kilogram.

OpenStudy (anonymous):

(a) 9.6 (b) 10.125 (c) 10.1 (d) 9.65 (e) 12.5 In kilograms

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

what is the volume of this cube?

OpenStudy (anonymous):

That is the whole question

jimthompson5910 (jim_thompson5910):

you use the info given to find the volume of the cube hint: V = s^3

jimthompson5910 (jim_thompson5910):

s = side length of the cube

OpenStudy (anonymous):

3.375

jimthompson5910 (jim_thompson5910):

now you'll multiply that by 3 since "Metal screws weigh 3 kilograms per cubic meter" then round that result to the nearest tenth

OpenStudy (anonymous):

Ty

jimthompson5910 (jim_thompson5910):

no problem

OpenStudy (anonymous):

In a seminar, the ratio of the men to women is 5:9. If there are 56 people in the seminar, how many are women?

jimthompson5910 (jim_thompson5910):

hints: m = # of men w = # of women " the ratio of the men to women is 5:9" so 5/9 = m/w there are 56 people total, so, m+w = 56. Solve for w to get w = 56-m you can plug that into the first equation to get this new equation below \[\Large \frac{5}{9} = \frac{m}{56-m}\]

OpenStudy (anonymous):

Dont get it

jimthompson5910 (jim_thompson5910):

further hint: cross multiply to get \[\Large \frac{5}{9} = \frac{m}{56-m}\] \[\Large 5(56-m) = 9m\] and solve for m

OpenStudy (anonymous):

Cant seem to find the answer :/

jimthompson5910 (jim_thompson5910):

give it another shot and post what you have so far in terms of your steps/work

jimthompson5910 (jim_thompson5910):

that way I can see where you're going wrong